A diprotic acid, H2A, has acid dissociation constants of Ka1=1.01×10−4 and Ka2=4.08×10−12. Calculate the pH and molar concentrations of H2A, HA−, and A2−at equilibrium for each of the solutions.

A diprotic acid, H, A, has acid dissociation constants of Kal = 1.01 x 104 and K22 = 4.08 x 10-12. Calculate the pH and molarA 0.176 M solution of NaHA. pH= pH= ] [H,A1 = C HA) = M [AP] = A 0.176 M solution of Na,A. pH= C HA] = M [AP]-C
Answer #1

* Сн, A) . 174 М — н, 4 HT (а. HT - HT (9) , Ко - }, о х 12 1 М. p. 1-4 2 3 - о. | 24 - и hty - 2 /o/x108. t= t - p. 124 г яГО. 12 п) - Na, HA] - Гні) hе - * * д”, кач, оr fo 11 р, ??-? 17 я в 1, ор х/, 12 д, 2) = - к Ко. », t, за - А - ---- limshi(o, 19cm] [4, 4) - (42) А - = AA + ой 7 Кое - . : І. р. 1 24 4.08 x 10th o. | 1 - 1 2 4.08 ? 2 чух 13. • 2 • ЧX tо ? О, 1-4